OpenStudy (anonymous):

What is the sum of the arithmetic sequence 8,14,20... , if there are 22 terms? answers to choose from. a)1,558 b)1,560 c)1,562 d)1,564

OpenStudy (anonymous):

@spectrum

OpenStudy (anonymous):

[8 + (0)(6)] + [8 + (1)(6)] + [8 + (2)(6)] + ... + [8 + (21)(6)] = (22)(8) + (6)(21)(22)/2 1 + 2 + 3 + ... + (n - 1) + n = n(n + 1)/2

OpenStudy (anonymous):

okay i believe the answer is 1,560 is this correct?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

well i wonder what i did wrong?? then it has to be 1,562 it is the closest to the answer i got right

OpenStudy (anonymous):

All the answers are 2 apart, so don't go by "closeness". Suppose you weren't given any choices and had nothing to choose from and had to get it by scratch. I already worked out the problem for you: (22)(8) + (6)(21)(22)/2 Just multiply it out. No guessing!

OpenStudy (anonymous):

i got 1,474 but that isnt any of the answers i posted:(

OpenStudy (anonymous):

If you are having problems with multiplying, I can't help you. Good luck.

OpenStudy (anonymous):

oh i see what i did wrong for got to divide @tcarroll010 thankyou

OpenStudy (kaylaprincess):

answer : 1562